A mass m, free to move along a line is attracted towards a given point on the line with a force proportional to its distance from the given point. If the mass starts from rest at a distance Xo from the given point, show that the mass moves in a simple harmonic motion.
1
Expert's answer
2019-03-29T11:16:29-0400

Let P - given point.

Coordinate x of P = 0.

Coordinate x of m = x.


Then "F = -kx" , so

"ma = -kx \\
ma + kx = 0 \\
m\ddot{x} + kx = 0"

We have linear differential equation,

Characteristic equation:

"m{\alpha}^2 + k = 0 \\
\alpha = \pm i \cdot \omega, 
 \omega = \sqrt{\frac{k}{m}}"

So

"x = C_1 \cdot e^{t\alpha_1} + C_2 \cdot e^{t\alpha_2} = C_1(e^{it\omega})+C_2(e^{-it\omega})"


with Euler's formula:

"x = C_1(\cos(\omega t) + i\sin(\omega t)) + C_2(\cos(-\omega t) + i\sin(-\omega t))"

"x = (C_1 + C_2)cos(\omega t)) + (iC_1 - iC_2)(sin(\omega t))"


Let:

"C_1 + C_2 = C_3 \\
iC_1 + iC_2 = C_4 \\
\sqrt{C_3^2+C_4^2} = A \\"

Then:

"(\frac{C_3}{A})^2 + (\frac{C_4}{A})^2 = 1 \\
\frac{C_3}{A} = \sin \phi \\
\frac{C_4}{A} = \cos \phi"


"x = C_3\cos(\omega t) + C_4\sin(\omega t)"

"x = A(\sin \phi \cos(\omega t) + \cos \phi \sin(\omega t))"


"x = A\sin(\omega t + \phi)"

- harmonic motion equation.